Circular No. 7614

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001az IN UGC 10483 T. Puckett and M. Peoples, Mountain Town, GA, report the discovery

of an apparent supernova (mag 17.1) on an unfiltered CCD frame (limiting mag 19.2) taken with the Puckett Observatory 0.30-m automated supernova patrol telescope on Apr. 27.39 UT. SN 2001az is located at α = 16^{h}34^{m}27^{s}.71, δ = +76^{o}01′46′′.8 (equinox 2000.0), which is 4′′.9 west and 0′′.3 north of the center of UGC 10483. The new object was also present on an unfiltered CCD frame taken on Apr. 28.16, but it does not appear on multiple unfiltered images taken by Puckett from 1999 Aug. to 2000 May, and it does not appear on Palomar Sky Survey images taken on 1994 June 2, 1992 Apr. 10 (limiting mag ∼ 21.0), and 1955 May 20 (limiting mag ∼ 19.7).

SUPERNOVA 2001ba IN MCG –05-28-1 R. Chassagne, Ste. Clotilde, Île de Réunion, reports the discovery of

an apparent supernova (mag ∼ 15.9) on unfiltered CCD images taken with a 0.3-m telescope on Apr. 27.8 and 28.7 UT. SN 2001ba is located at α = 11^{h}38^{m}01^{s}.76, δ = −32^{o}19′51′′.0 (equinox 2000.0), which is 18′′.7 east and 21′′.8 south of the nucleus of MCG –05-28-1 (= ESO 439-G18). An image of the same field taken on Feb. 19.9 showed nothing at this position (limiting mag ∼ 18).

P. Nugent and L. Wang, Lawrence Berkeley National Laboratory, report: “A spectrum (range 400–800 nm) obtained on Apr. 30.17 UT at the Kitt Peak 2.1-m telescope (+ GCAM) shows that SN 2001ba is a type-Ia supernova near maximum light. Adopting the NASA/IPAC Extragalactic Database recession velocity of 8819 km/s for the host, we find that the minimum of the Si II feature is at a velocity of 10 900 km/s.”

SUPERNOVA 2001bb IN IC 4319 M. Schwartz, Cottage Grove, OR; and W. D. Li, University of Califor-

nia at Berkeley, on behalf of LOTOSS (cf. IAUC 7514), report the discovery of an apparent supernova on unfiltered images taken with the 0.5-m Tenagra III automated telescope on Apr. 29.3 (mag ∼ 16.5) and 30.2 UT (mag ∼ 16.3). SN 2001bb is located at α = 13^{h}43^{m}25^{s}.31, δ = −29^{o}48′13′′.1 (equinox 2000.0), which is ∼ 16′′.8 west and 1′′.4 south of the nucleus of IC 4319. An image taken by the Katzman Automatic Imaging Telescope on Apr. 15.3 showed nothing at this position (limiting mag ∼ 18.5).

2001 April 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green